OpenStudy (anonymous):

4. Given the linear equation y = 5x – 1, find the y-coordinates of the points (-1, ), (2, ), and (1, ). Please show all of your work. Plot these points and graph the linear equation

OpenStudy (zepp):

Simply replace x in your equation by the number they give you, and find the corresponding y value for the x value.

OpenStudy (zepp):

Let me do the first one for you; We are looking for the corresponding y-value for -1 Replace x by -1 y = 5(-1) – 1 y = -5 -1 y = -6 The corresponding y value for x (-1) would be -6; (-1,-6)